OpenStudy (anonymous):

find all the roots of the polynomial 3x^4-17x^3+21x^2+3x-2. express your answer in simplest radical form. please help! x

OpenStudy (anonymous):

please help!

OpenStudy (rajee_sam):

first root you have to find by trial and error. Look for the factors of 2 (Your constant term). +1, -1, +2, -2 are those. Now look for the roots of 3 (your leading coefficient) , +1, -1, +3, -3 Now your possible real roots are +1, -1, +2, -2, +1/3, -1/3, +2/3, -2/3 This I got by dividing all the roots of 2 by all the roots of 3 that I listed above. Now you have to substitute each value in your function to see which turns the value of the function you Y = 0. That will be one root. So lets try

OpenStudy (anonymous):

3x^4-17x^3+21x^2+3x-2 =3x^4-6x^3-11x^3+22x^2-x^2+2x +x-2 = 3x^3(x-2) -11x^2 (x-2) -x(x-2)+1 (x-2) =(x-2)(3x^3-11x^2-x+1)

OpenStudy (luigi0210):

Find rational roots using p/q, one of the roots is -1/3

OpenStudy (luigi0210):